OpenStudy (anonymous):

A box contains the letters M I S S I S S I P P I. What is the probability of selecting an S first and a P second, if: a. the first letter is not replaced? b. the first letter is replaced.

OpenStudy (anonymous):

4/11*2/10

OpenStudy (anonymous):

because there are 4 S out of 11 letters, there is 2 P out of the remaining 10 letters

OpenStudy (anonymous):

so the a. the first letter is not replaced what do i put?

OpenStudy (anonymous):

thus the number in the denominator goes from 11, the original number o letters in the word to 10, the remaining letters you can choose from

OpenStudy (anonymous):

so what do i write????>?>

OpenStudy (anonymous):

4/11*2/10=8/110 which reduces to 4/55

OpenStudy (anonymous):

so that with the first letter not being replaced?

OpenStudy (anonymous):

yes, you leave the S out when you get ready to draw the 2nd time

OpenStudy (anonymous):

so the 2nd time on b. is the first letter is replaced what do i right

OpenStudy (anonymous):

then you take 4/11*2/11 since there is 11 letters on each draw

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

if the 1st letter is replaced the answer is 4/11*2/11 = 8/121

OpenStudy (anonymous):

Lols i really don't get this but ok can u explain clearly more about how u got all those Numbers